a perimeter of a rectangle is 78 feet. its length is seven feet more than its width. what are the dimensions of the rectangle
Answer by sachi(548)

let the width is w
then length is w+7
so perimeter=2(w+7+w)=78
or 2w+7=78/2=39
or 2w=39-7=32
or w=32/2=16 & l=16+7=23
ans